OpenStudy (zzr0ck3r):

Fun challenge. Knowing only that between any two real numbers is a rational number. Show that between any two real numbers is an irrational number.

OpenStudy (perl):

\(\large a+\frac{\sqrt{2}(b-a)}{2}\) is between \( a\) and \( b\) and is irrational

OpenStudy (rsadhvika):

Consider any two real numbers \(r_1\) and \(r_2\) we know that there exists a rational number \(a\) in \((\pi+r_1, \pi + r_2)\). Then \(a-\pi \in (r_1,r_2)\).

OpenStudy (zzr0ck3r):

this needs pi is irrational, which is not so trivial. But sqrt(2) is...

OpenStudy (rsadhvika):

Ahh right, any irrational number will do !

OpenStudy (zzr0ck3r):

Also needs rational + irrational is irrational

OpenStudy (zzr0ck3r):

That is how I answered it today at our tudor table, but I really liked the question and thought it was very cute

OpenStudy (zzr0ck3r):

outside of the famous proofs, I think that is one of my favs (for the little guys). I also like that identities are unique proof. \(e_1=e_1e_2=e_2\) very cute as well.

OpenStudy (perl):

identity for a group?

OpenStudy (zzr0ck3r):

yeah or monoid or ....

OpenStudy (zzr0ck3r):

You technically won, but I was not gonna medal a 99er

OpenStudy (zzr0ck3r):

:)

OpenStudy (zzr0ck3r):

Would love to see other ideas also.

OpenStudy (perl):

rsadvika's proof is neat ;)

OpenStudy (across):

Here's a lengthier proof: Let \(r_1\) and \(r_2\) be any two different real numbers, let \(q\) be a rational number between \(r_1\) and \(r_2\), and let \(k\) be an irrational number. Since both the sum and the product of a rational number and an irrational number is an irrational number, it follows that \(\left\{q+k/n\right\}_{n=1}^\infty\) is a sequence of irrational numbers converging to \(q\). This proves that there are infinitely-many irrational numbers between \(r_1\) and \(r_2\).

OpenStudy (across):

I should say countably infinite...

OpenStudy (zzr0ck3r):

way to much analysis :) wth does converge mean

OpenStudy (across):

It just means that\[\lim_{n\to\infty}q+\frac kn=q.\]

OpenStudy (across):

@perl's proof breaks down for \(a=0\) and \(b=\sqrt2\).

OpenStudy (zzr0ck3r):

I was joking about the sequence thing. I was more pointing out the tools needed for that proof.

OpenStudy (zzr0ck3r):

and rsadhvika gives us density also, because we just do it countably infinitely many times:)

OpenStudy (zzr0ck3r):

but I like it. Keep em coming !!!

OpenStudy (across):

You could also say that if there are no irrational numbers in \(\left[r_1,r_2\right]\), then it is countably infinite, which is a contradiction.

OpenStudy (zzr0ck3r):

but you need cantor for the contradiction. love me some cantor

OpenStudy (across):

I know, I know, xD.

OpenStudy (zzr0ck3r):

you are way to quick though :)

OpenStudy (zzr0ck3r):

Cantor might be to much for most of the people here, but I bet the enumeration of the rationals would not be. The zig zag thingy

OpenStudy (kainui):

If you choose the arbitrary real numbers \(a=(2+\sqrt{2})c\) and \(b=\sqrt{2}d\) then perl's condition: \[a + \frac{\sqrt{2}(b-a)}{2} =c+d \] So in general there are infinitely many holes in that solution. It looks like you might accidentally end up with a limit if you're not careful with this kind of reasoning though if you suggest wiggling in the bounds to some other number, what's to guarantee that next set of numbers to also not be a miss and accidentally fall into an infinite regress?